# SAGE: Streaming Agreement-Driven Gradient Sketches for Representative Subset Selection

### v1

@@ 4 Related Work
Coreset selection and data pruning for efficient training have been explored via submodular coverage, gradient matching, and influence functions, among other ideas [12, 9, 7, 8, 10, 11]. CRAIG and GLISTER exemplify selection by coverage and generalization proxies, respectively, while GradMatch formulates an explicit gradient matching objective that scales quadratically in the number of examples [9, 8, 7]. More recent directions examine distributionally robust pruning and stepped greedy strategies aimed at favorable empirical trade-offs [22, 23, 24]. However, many of these approaches either incur $\Theta(N^{2})$ pairwise computations or require storing/operating on an explicit $N\times D$ gradient matrix, and several rely on bilevel or proxy objectives whose stability can vary across datasets.

### v2

@@ 4 Related Work
Coreset selection and data pruning for efficient training have been explored via submodular coverage, gradient matching, and influence functions, among other ideas [12, 9, 7, 8, 10, 11]. Craig and Glister exemplify selection by coverage and generalization proxies, respectively, while GradMatch formulates an explicit gradient-matching objective that scales quadratically in the number of examples [9, 8, 7]. More recent directions examine distributionally robust pruning and stepped greedy strategies aimed at favorable empirical trade-offs [22, 23, 24]. Complementary to these, Graft employs Fast MaxVol sampling on low-rank feature projections with dynamic gradient-alignment adjustments to perform in-training subset selection [25], while Sage uses a Frequent Directions sketch with gradient-agreement scoring to perform streaming selection in constant memory and two passes, avoiding explicit $N^{2}$ similarities [25]. In parallel, Market-Based Data Subset Selection casts multi-criteria example utility into a convex cost/pricing framework, enabling principled aggregation of heterogeneous signals and tunable trade-offs [26]. However, many existing approaches either incur $\Theta(N^{2})$ pairwise computations or require storing/operating on an explicit $N\times D$ gradient matrix, and several rely on bilevel or proxy objectives whose stability can vary across datasets.